OpenStudy (anonymous):

F(x)=f(f(x)) and G(x)=(F(x))2 and suppose that f(5)=11,f(11)=2,f′(11)=6,f′(5)=5. What is G'(5) ?

OpenStudy (anonymous):

i got F'(5) to be 30 but i'm not getting the right answer for G'(5)

OpenStudy (freckles):

let me check your first one... \[F(x)=f(f(x)) \\ F'(x)=f'(x) \cdot f'(f(x)) \\ \text{ now replace } x \text{ with 5} \text{ \to find } F'(5) \\ F'(5)=f'(5) \cdot f'(f(5))=5 \cdot f'(11)=5 \cdot 6=30\] first one looks incredible

OpenStudy (freckles):

what is G(x)?

OpenStudy (freckles):

\[G(x)=(F(x))^2\]

OpenStudy (freckles):

is that right?

OpenStudy (freckles):

\[G'(x)=2 (F(x))^{2-1} \cdot (F(x))'\] and you already found F'(x) so replace it down here

OpenStudy (anonymous):

okay so i got 2(f(f(5))*30 and then 2f(11) * 30 and then 2* 2* 30 and then 120

OpenStudy (freckles):

\[F(x)=f(f(x)) \\ F(5)=f(f(5))=f(11)=2 \\ \\ G'(5)=2F(5)F'(5)=2(2)(30)\] had to do myself to make sure yes that looks good :)

OpenStudy (anonymous):

ok thanks